### v1

# Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems

Abstract.
Large language model-based agents are increasingly used in recommender systems (Agent4RSs) to achieve personalized behavior modeling. Specifically, Agent4RSs introduces memory mechanisms that enable the agents to autonomously learn and self-evolve from real-world interactions. However, to the best of our knowledge, how robust Agent4RSs are remains unexplored. As such, in this paper, we propose the first work to attack Agent4RSs by perturbing agents’ memories, not only to uncover their limitations but also to enhance their security and robustness, ensuring the development of safer and more reliable AI agents.
Given the security and privacy concerns, it is more practical to launch attacks under a black-box setting, where the accurate knowledge of the victim models cannot be easily obtained. Moreover, the practical attacks are often stealthy to maximize the impact. To this end, we propose a novel practical attack framework named DrunkAgent. DrunkAgent consists of a generation module, a strategy module, and a surrogate module. The generation module aims to produce effective and coherent adversarial textual triggers, which can be used to achieve attack objectives such as promoting the target items. The strategy module is designed to ‘get the target agents drunk’ so that their memories cannot be effectively updated during the interaction process. As such, the triggers can play the best role. Both of the modules are optimized on the surrogate module to improve the transferability and imperceptibility of the attacks. By identifying and analyzing the vulnerabilities, our work provides critical insights that pave the way for building safer and more resilient Agent4RSs. Extensive experiments across various real-world datasets demonstrate the effectiveness of DrunkAgent.
Adversarial Attacks, Recommender Systems, Large Language Model, Generative Agents, Collaborative Filtering
copyright: noneccs: Security and privacy Web application security
1. Introduction
Recommender systems (RSs) have evolved into a vital component of the modern digital ecosystem. From e-commerce sites (e.g., Amazon, Taobao) and streaming platforms (e.g., YouTube, TikTok) to social media networks (e.g., Facebook, Instagram), RSs power personalized experience, drive user engagement of these online platforms, and further influence market dynamics through item promotions (Yang et al., 2024, 2023). The foundation of the RSs lies in modeling users’ preferences based on their historical interactions with items, such as purchases and ratings, known as collaborative filtering (CF). Despite the success of CF techniques (Koren et al., 2009; He et al., 2017, 2020), they often struggle with data sparsity and cold-start problems (Sachdeva and McAuley, 2020). To alleviate the problems, the auxiliary information with rich semantics (e.g., item descriptions and user reviews) has been integrated into RSs (Zhang et al., 2019). Given the increasing volume and complexity of recommendation data, the RSs encounter difficulties in scaling effectively and capturing textual knowledge about users and items sufficiently, due to they generally rely on traditional deep learning (DL) models like CNNs and RNNs (Zhao et al., 2023).
Figure 1. An exemplar of attacks on agent-based recommender systems, where the target item’s descriptions have been manipulated to perturb the memory of the target agent for promoting the target item to more normal users. By investigating this, our work reveals emerging vulnerabilities in memory-augmented agents and provides insights for building more secure recommender systems.
Recently, large language models (LLMs) have shown potential to mitigate the above problems and revolutionize RSs (Zhao et al., 2023), as they are primarily built upon advanced transformer-based architectures and pretrained on diverse and extensive datasets. Most of works carefully design textual prompts for LLMs based on user-item interactions (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Wu et al., 2024). As such, the abundant knowledge inside the LLM’s training corpora can be fully exploited. However, the LLM-based RSs (LLM4RSs) struggle to capture underlying behavioral patterns, due to the gap between universal language modeling and personalized behavior modeling (Zhang et al., 2024b). Integrating agents built upon LLMs equipped with additional modules offers a promising solution (Durante et al., 2024; Huang et al., 2024), allowing to autonomously explore and learn from external sources. Among all the added modules, memory is a key component to support agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task. More broadly, environment can be any contextual factors (e.g., item titles, user-item interactive behaviors) that influence the agent’s decisions (Zhang et al., 2024a). interactions, which plays a crucial role in helping agents accumulate historical experience, retrieve informative knowledge and self-evolve in real-world environments (Zhang et al., 2024a; Wang et al., 2024b). Consequently, the agent-based RSs with powerful LLMs (Huang et al., 2024; Zhang et al., 2024b) can also leverage memory modules that store the information explored from the external environments to further improve the recommendation quality. Fig.  1 gives an example of an item agent interacting with its situated environment (e.g., users and other items), where the learning knowledge will be placed in its memory. Despite the aforementioned success, the safety vulnerability of the generative agent-based RSs (Agent4RSs) under adversarial attacks has yet to be explored. Such attacks can gain insights into the robustness of the systems and help develop effective defensive mechanisms. As recommender agents are integrated into more complex real-world environments, understanding how these attacks might compromise their behaviors is critical for building trustworthy and robust recommender agents (Chua et al., 2024).
Due to security and privacy considerations, the access of accurate knowledge of the victim model, including architectures, gradients, parameters, etc., is generally restricted in practice. As such, practical attacks are mainly under a black-box setting (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023). Practical attacks are often required to have transferability (i.e., effective against different black-box RSs). Moreover, to maximize the impact of the attack and the number of affected users, practical attacks are often rewarded by imperceptibility (i.e., difficult to be discovered) (Liu and Larson, 2021).
Conventional black-box shilling attacks (Wang et al., 2024a), which inject well-crafted fake user profiles (e.g., a set of users’ numerical ratings on items) into the victim model for certain attack objectives (e.g., promote their target items for profits), have been shown to effectively poison DL-based RSs. However, most attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses (Ning et al., 2024). Item providers are generally responsible for creating and maintaining their own product representations on the platforms with recommendation services in practice (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021) (e.g., YouTube allows creators to upload item titles and Amazon enables sellers to upload item descriptions), and current RSs depend heavily on these representations (Zhang et al., 2024d; Zhao et al., 2023). As such, security vulnerabilities in these systems naturally emerge. The vulnerabilities can easily be exploited by profit-driven merchants manipulating product representations, as shown in Fig.  1. Compared to the shilling attacks that involve hiring online writers to introduce fake reviews to indirectly promote the items, this type of attack is more direct, efficient, and lower in economic cost (Zhang et al., 2024d).
Recent black-box textual attacks on LLM4RSs are far from practical. Classical text attacks that are successful in text mis-classifications, such as TextFooler (Jin et al., 2020) and PuncAttack (Formento et al., 2023), have been directly applied for LLM4RSs (Zhang et al., 2024d). However, the effectiveness of the attacks is limited, due to there is a gap between the attack goals of the text attacks and RS attacks. Trivial and ChatGPT attacks (Zhang et al., 2024d), on the other hand, are proposed to be tailored for LLM4RSs, which increases the items’ attractiveness by appending positive words to the titles. Nevertheless, such attacks are not optimized on any black-box RSs, making them poorly transferable to the Agent4RSs. Moreover, since the attacks mainly target modifications to item titles, where titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
To tackle the above problems, in this paper, we propose a new practical attack framework named DrunkAgent. Instead of manipulating the item titles, to enhance the imperceptibility of attacks, we propose to perturb the description of the target item, where length and context can easily hide changes. DrunkAgent is composed of three modules: generation module, strategy module and surrogate module. The generation module is designed to produce effective adversarial textual description triggers, which can be used to promote the target items to as many normal users as possible on the agent-powered RSs. The triggers are coherent, semantically meaningful and human-readable text, making the attack stealthy. The strategy module aims to ‘get the target agents drunk’, that is, to confuse the agents’ memory mechanisms so that they cannot be effectively updated during agent-environment interactions. In this way, the triggers can play the best role. To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the generation module and the strategy module are optimized on the surrogate module to improve the transferability and imperceptibility of the practical attacks.
Beyond demonstrating the attack capabilities of DrunkAgent, our broader goal is to reveal and understand the overlooked vulnerabilities in Agent4RSs. Specifically, by analyzing how agents’ memory mechanisms can be influenced through carefully crafted and stealthy textual triggers, our work provides a new lens into the emerging attack surfaces introduced by memory-augmented, LLM-based agents. We believe that this exploration can offer valuable insights to the community for designing more resilient agent architectures, guiding future efforts on defense mechanisms such as memory validation or input sanitization.
Our main contributions are summarized as follows:
  • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs. To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agent-powered RSs, aiming to inform future robustness research.
  • We propose to perturb the description of the target item for effective memory perturbations, which is an unnoticeable and crucial property of current agentic recommendations.
  • We propose a novel practical attack framework on the agent-powered RSs named DrunkAgent, which can create adversarial descriptions of target items with transferability and imperceptibility.
  • Our DrunkAgent significantly outperforms state-of-the-art black-box attacks against various LLM-based agentic RSs on real-world datasets in terms of transferability and imperceptibility. Moreover, state-of-the-art defense strategies are ineffective to DrunkAgent, demonstrating the need for improved defense mechanisms to guide more secure Agent4RSs.
2. Problem Formulation
2.1. Victim LLM-powered Agentic RSs
We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively. $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile). LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (Geng et al., 2022; Zhao et al., 2023; Bao et al., 2023; Zhang et al., 2023). Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
(1) $\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$
where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$. $\oplus$ denotes the textual string concatenation. $\mathcal{P}$ stands for prompt templates of recommendations. $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items. $\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (Xu et al., 2021). $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$K$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions. In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents. Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
2.2. Threat Model
2.2.1. Attacker’s Knowledge:
As the works (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023) point out, the inherent details (e.g., the architectures, gradients and parameters) of the victim RSs often cannot be obtained in practice. As such, our attack is performed under the black-box settings. In addition, the attacker is only able to access publicly available recommendation data (e.g., item titles and user reviews that can be collected directly by browsing or writing crawlers), due to the openness nature of RSs (Si and Li, 2020; Yang et al., 2023).
2.2.2. Attacker’s Capability:
Unlike another state-of-the-art LLM-powered RS attack (Ning et al., 2024) with an unrealistic assumption of $\mathcal{P}$ and $\mathcal{V}_{u}$ can be easily perturbed (that will be detailed in Section 5), we consider that the attackers have limited perturbation capability in practice. Specifically, they can only interfere with the representation of the target item, which is feasible and can be easily accomplished via a dedicated API, as the existing practical attacks do (Zhang et al., 2024d; Yang et al., 2024; Liu and Larson, 2021).
2.2.3. Attacker’s Objectives:
As mentioned in Section 1, two objectives: transferability and imperceptibility are considered for the practical attack. Given that push attacks are generally used by unscrupulous parties on RSs for profits in practice (Si and Li, 2020) with the aim of promoting target items to as many normal users as possible (Zhang et al., 2021), we thus focus on such attacks in this paper, as normally do in recent works (Wang et al., 2024a; Zhang et al., 2024d). Our first goal is to maximize the hit ratio of the target item $t$ on a surrogate RS (denoted as $\text{Agent}_{\Phi}$ with parameters $\Phi$) that is different from the victim black-box RS, ensuring transferability. Our second goal is to make the descriptions of $t$ as stealthy as possible, achieving imperceptibility. As such, the surrogate RS is used to optimize attacks to avoid frequently querying the victim models to induce suspicion.
Figure 2. Practical Attack Framework DrunkAgent Overview
3. DrunkAgent
In this section, we present a practical attack framework, DrunkAgent, for gaining insights into the robustness of autonomous agent-based RSs with powerful LLMs and related defense models in practice. DrunkAgent is composed of three attack modules. Fig.  2 gives an overview of DrunkAgent, where each component will be described in the following.
3.1. Surrogate Module
We leverage a surrogate RS $\Phi$ to measure how effective the attack is under the black-box settings, as inspired by the existing works (Lin et al., 2022; Tang et al., 2020; Yang et al., 2023, 2024). It is used to optimize the generation module and the strategy module, making the attack transferable and imperceptible.
Since our attack targets SOTA LLM-powered RSs, we adopt powerful LLMs as recommendation agents. Given the transparency, flexibility and scalability of the open-source LLM models, they can be considered for building this module. Moreover, due to our attack focus on adversarial text description generation tasks, we consider an autoregressive language model trained with a causal language modeling objective (Radford et al., 2019; Brown et al., 2020). There are some notable models, such as LLaMA (Touvron et al., 2023a), LLaMA 2 (Touvron et al., 2023b) and LLaMA 3 (Dubey et al., 2024), BLOOM 176B (Le Scao et al., 2023) and Falcon 180B (Almazrouei et al., 2023). Currently, Meta’s LLaMA family has become the most influential open LLM and has been widely adopted in the recent recommendation community (Zhao et al., 2023; Bao et al., 2023). To ensure the effectiveness of attacks, we apply the latest generative model: LLaMA 3, so that valuable feedback can be provided even with limited available resources. Specifically, LLaMA 3 uses an optimized transformer architecture (Vaswani et al., 2017). Its instruction-tuned version uses supervised fine-tuning and reinforcement learning from human feedback (RLHF) to align with human preferences for helpfulness and safety (AI, 2024), as shown in Fig.  2.
In addition to conducting a strict selection of recommendation agents, we develop the two modules for effective target item’s adversarial description generation.
3.2. Generation Module
In this section, we aim to generate adversarial description triggers on the agent-powered RSs, which can promote the target items to as many normal users as possible. The triggers should have three critical properties: 1) Universality on any users and items; 2) Transferability among different black-box victim models; and 3) Imperceptibility to not easier to be discovered. To this end, we design sub-modules, as shown in Fig.  2 and described as follows.
3.2.1. Adversarial Prompting Engineering
Recently, many of the mainstream LLMs that have been widely used in the agentic recommendation community (Huang et al., 2024) (e.g., GPT-3.5-turbo, ChatGPT) are closed source, where the models’ underlying codes, weights, or other critical components are not released to the public and the models can only be invoked through their APIs. As a result, the agent-powered RSs generally adopt in-context learning (ICL) (Dong et al., 2022) to sufficiently take advantage of the generalization capability of LLMs on unknown tasks. To this end, this module carefully crafts prompts $\widetilde{\mathcal{X}_{t}}$ within the ICL framework to improve the quality of adversarial trigger generation.
Basic Prompt Template. We meticulously prepare prompt templates $\widetilde{\mathcal{P}}$ to provide context, goals, instructions and constraints so that the model knows ‘what needs to be done’ and ‘how to respond to’. The main components are described in the following.
  1. (1)
    Style Definition $\widetilde{p_{s}}$. There are two common styles of prompts: task-oriented and role-oriented (Zhu et al., 2023a). To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig.  2.
  2. (2)
    Task Goal $\widetilde{p_{g}}$. Since a recommender system is essentially a ranker (Zhang et al., 2019), we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
  3. (3)
    Recommendation Instruction $\widetilde{p_{i}}$. We introduce highly detailed instructions to guide the model to perform accurate recommendations. For example, ‘Please sort CDs by considering the correlation between your preferen -ces/dislikes and the features of the CDs. The higher the CD rank, the more you like it.’ in the shown example.
  4. (4)
[…]
Hence, the adversarial prompt template can be formulated as:
(2) $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
To increase the attack effectiveness, we introduce position bias (Zhang et al., 2024b) into $\widetilde{\mathcal{P}}$, where the other candidates items are placed before the target item, as shown in the example of Fig.  2.
General User Memories. It is difficult to obtain the optimized memories of user agents (i.e., language-based parametric embeddings of the agentic RSs (Zhang et al., 2024b)), due to the practical attack is launched under the black-box settings. Moreover, the adversarial description triggers are designed to be effective for general users, rather than specific user groups (e.g., users who are only obsessed with Progressive Metal music), so that the target items can be promoted to as many users as possible. Therefore, to enhance the universality of the triggers on users, we use primitive and general user descriptions as user memories $\widetilde{m_{u}}$, such as ‘I enjoy listening to CDs very much.’ in the given example, as the default user memories in the existing works (Zhang et al., 2024b)(Huang et al., 2024).
Popular Item Memories. Given that popular items usually have a high probability of appearing at the top of the users’ recommendation lists, Bandwagon attack (Si and Li, 2020) introduces popular items into fake user profiles to poison traditional RSs for increasing the popularity of the target item and Expert attack on visually-aware RSs (Liu and Larson, 2021) creates adversarial images by decreasing the representation distance to popular items for improving the exposure rate of the target item. Inspired by this, we select popular items as candidate items for $\widetilde{\mathcal{X}_{t}}$ to introduce popularity bias (Zhang et al., 2024b) to further enhance the effectiveness of adversarial description triggers. The basic idea is that the target items should be promoted to more normal users if they can be ranked ahead of the popular items. Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items,
(3) $\mathcal{M}_{p}=\{m_{v}\},$
where $m_{v}$ is the features of the item $v$ including titles and descriptions. $\{v\}_{v\in\mathcal{V}}$ is a sequence of popular item IDs, which is corresponding to the subscript of the item popularity sequence in descending order $\{|\mathcal{U}_{v}|\}$. The item popularity is item frequency here, due to this work focuses on an implicit task.
The descriptions of candidate items will greatly influence the feedback of the surrogate model. As with optimized user agent memories, the adversary is hard to access to internal memories of item agents in practice and is restricted to leverage publicly available data (e.g., item titles and descriptions). However, the raw descriptions (Ni et al., 2019) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations. Similar to the use of user memories, we can consider the default item memories for popular items. Nevertheless, the default ones generally only include identity information such as titles and categories (Zhang et al., 2024b), as the example “The CD is called ‘Brainwashed’. The category of this CD is: ‘Classic Rock; Album-Oriented Rock (AOR)’.” shows, which are overly simple and information-less, leading to the target items easier to be ranked at the top and the sub-optimal results will be produced. To tackle these issues, we propose to adopt a LLM with frozen parameters $\Psi$ to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations. Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items,
(4) $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p}),$
where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition and OpenAI’s GPT-4 (Achiam et al., 2023) is adopted to implement the LLM, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities. An example of producing a popular item memory is given in Fig.  2. To sum up, the adversarial prompt can be formulated as
(5) $\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus\widetilde{m_{t}},$
where $\widetilde{m_{t}}$ is the features of the target item. In addition to the fixed title, the adversarial description trigger (i.e., the memory of the item agent) will be optimized in the next sub-section.
Note that we give lightweight and efficient adversarial prompts for the surrogate RS that are different from the real victim RSs’. To ensure high-quality recommendations, more powerful and context-rich (e.g., retrieval-augmented) prompts will be used in the victims, which will be described in the evaluation section.
3.2.2. Greedy Search Algorithm
We propose a greedy search algorithm to optimize the adversarial description triggers, as inspired by the successes of the current jailbreak attacks on LLMs (Zou et al., 2023; Liu et al., 2024b). However, there are some essential differences (e.g., attack goals and victim models) between jailbreak attacks and our attacks. Jailbreak attacks are designed to undermine the LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection, whereas our attacks aim to promote the target items on the RSs powered by autonomous agents.
Adversarial Trigger Initialization. Given that LLMs should understand LLM-powered agents better, we leverage a LLM to generate some description candidates. To narrow search space by a large margin to increase the efficiency and effectiveness of the algorithm, the basic attack goals are incorporated into the prompts. In addition, the diversity and the generalizability are introduced to the prompts to help search for the global optimal solution and improve their universality on items, respectively. Formally,
(6) ~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus\widetilde{m_{t}}),$
where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM and the state-of-the-art GPT-4 is used to implement the LLM. $\widetilde{\mathcal{X}_{c}}$ denotes a sequence of prompt templates. An example of the initialization is given in Fig.  2.
Adversarial Trigger Optimization. The algorithm takes $\mathcal{E}$ epochs to obtain the optimal trigger, where the output candidates of each epoch will be the inputs of the next epoch for continuous optimization until the attack objective is achieved or the termination condition is reached (e.g., the maximum number of epochs is exhausted). To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the optimization is built on the surrogate module, as shown in Fig.  2.
For each epoch $k$, the algorithm first calculate the performance score for each input candidate $\widetilde{m}_{t_{i}}^{(k-1)}$, where $n$ well-performing trigger candidates will be maintained directly for the next optimization to ensure the stability. The higher the score, the more likely the candidate is to be kept. The score is calculated by the negative of the loss, i.e., $s_{i}^{(k-1)}=-\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}$. The smaller the loss, the more likely the target item is to be ranked high and the more effective the candidate is. We use the auto-regressive language generation loss (i.e., negative log-likelihood) to evaluate the discrepancy between the predictions and the target output $\widetilde{\mathcal{R}_{t}}$,
(7) $\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$
[…]
(8) $\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}))=\prod_{l=1}^{L}\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{r_{t_{l}}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}\oplus\widetilde{r_{t_{<l}}})),$
Then, the algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection, where the candidates with higher scores will have a greater probability of being selected into this subset. The selection probability of each candidate is calculated by a softmax function:
(9) $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}.$
For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts. For example, for the pairwise $<\widetilde{m}_{t_{i}}^{(k-1)},\widetilde{m}_{t_{j}}^{(k-1)}>$, $\widetilde{m}_{t_{i}}^{(k-1)}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$, the new candidates could be $\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)}]$. As a result, the new combinations of different words, phrases, and sentences are gained, which increases the feature diversity of candidates through extensive exploration, contributing to search for the optimal trigger.
Table 1. Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets. We use bold fonts to denote the best performance. Note that the attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
Victim RS AgentCF
Attack CDs & Vinyl Office Products Musical Instruments
H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3
Benign 0.0505 0.1111 0.1616 0.0505 0.0887 0.1140 0.0306 0.0612 0.0918 0.0306 0.0499 0.0652 0.0412 0.0515 0.1134 0.0412 0.0477 0.0787
DeepwordBug 0.0808 0.1919 0.3232 0.0808 0.1509 0.2166 0.0000 0.0306 0.0510 0.0000 0.0193 0.0295 0.0206 0.0515 0.1031 0.0206 0.0401 0.0659
PuncAttack 0.0505 0.1313 0.2323 0.0505 0.1015 0.1520 0.0612 0.1224 0.1531 0.0612 0.0999 0.1152 0.0103 0.0309 0.0515 0.0103 0.0233 0.0336
TextFooler 0.0707 0.1717 0.2626 0.0707 0.1344 0.1799 0.0510 0.0816 0.1224 0.0510 0.0703 0.0907 0.0206 0.0309 0.0515 0.0206 0.0271 0.0374
BertAttack 0.0303 0.1111 0.2020 0.0303 0.0813 0.1267 0.0000 0.0306 0.0510 0.0000 0.0193 0.0295 0.0309 0.0515 0.0825 0.0309 0.0439 0.0594
TrivialInsertion 0.0202 0.1111 0.1818 0.0202 0.0776 0.1129 0.0306 0.0816 0.1531 0.0306 0.0628 0.0985 0.0103 0.0206 0.0619 0.0103 0.0168 0.0374
ChatGPTAttack 0.0606 0.1515 0.2626 0.0606 0.1180 0.1735 0.0510 0.1122 0.1633 0.0510 0.0896 0.1152 0.0103 0.0412 0.0825 0.0103 0.0298 0.0504
DrunkAgent 0.4040 0.4141 0.4343 0.4040 0.4104 0.4205 0.2449 0.2449 0.2551 0.2449 0.2449 0.2500 0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
Victim RS AgentRAG
Attack CDs & Vinyl Office Products Musical Instruments
H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3
Benign 0.0505 0.1111 0.2222 0.0505 0.0887 0.1443 0.0510 0.0510 0.0816 0.0510 0.0510 0.0663 0.0206 0.0412 0.1031 0.0206 0.0336 0.0646
DeepwordBug 0.0202 0.1010 0.2626 0.0202 0.0712 0.1520 0.0000 0.0204 0.0816 0.0000 0.0129 0.0435 0.0000 0.0206 0.0722 0.0000 0.0130 0.0388
PuncAttack 0.0404 0.1818 0.2626 0.0404 0.1296 0.1700 0.0510 0.0612 0.1122 0.0510 0.0575 0.0830 0.0206 0.0309 0.0619 0.0206 0.0271 0.0426
TextFooler 0.0606 0.1111 0.2222 0.0606 0.0925 0.1480 0.0510 0.1122 0.1429 0.0510 0.0896 0.1050 0.0000 0.0103 0.0412 0.0000 0.0065 0.0220
BertAttack 0.0404 0.1010 0.2020 0.0404 0.0786 0.1291 0.0000 0.0000 0.0102 0.0000 0.0000 0.0051 0.0103 0.0309 0.0309 0.0103 0.0233 0.0233
TrivialInsertion 0.0000 0.0707 0.2020 0.0000 0.0446 0.1103 0.0408 0.0816 0.1327 0.0408 0.0666 0.0921 0.0103 0.0206 0.0412 0.0103 0.0168 0.0271
ChatGPTAttack 0.0606 0.1616 0.2525 0.0606 0.1243 0.1698 0.0816 0.1122 0.1735 0.0816 0.1009 0.1316 0.0103 0.0103 0.0412 0.0103 0.0103 0.0258
DrunkAgent 0.3131 0.3232 0.3333 0.3131 0.3195 0.3246 0.1837 0.1837 0.2143 0.1837 0.1837 0.1990 0.1340 0.1443 0.1443 0.1340 0.1405 0.1405
Victim RS AgentSEQ
Attack CDs & Vinyl Office Products Musical Instruments
H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3
Benign 0.0404 0.1111 0.1515 0.0404 0.0850 0.1052 0.0306 0.0408 0.0408 0.0306 0.0371 0.0371 0.0103 0.0206 0.1134 0.0103 0.0168 0.0632
DeepwordBug 0.1111 0.2323 0.3232 0.1111 0.1876 0.2330 0.0408 0.0510 0.0714 0.0408 0.0473 0.0575 0.0206 0.0309 0.0928 0.0206 0.0271 0.0581
PuncAttack 0.1212 0.1818 0.3030 0.1212 0.1595 0.2201 0.1224 0.1327 0.1327 0.1224 0.1289 0.1289 0.0515 0.0619 0.1340 0.0515 0.0581 0.0941
TextFooler 0.1616 0.1818 0.2121 0.1616 0.1744 0.1895 0.0408 0.0612 0.0816 0.0408 0.0537 0.0639 0.0103 0.0206 0.0515 0.0103 0.0168 0.0323
BertAttack 0.0707 0.0909 0.1919 0.0707 0.0835 0.1340 0.0102 0.0102 0.0306 0.0102 0.0102 0.0204 0.0206 0.0515 0.1340 0.0206 0.0401 0.0814
TrivialInsertion 0.0202 0.0808 0.2727 0.0202 0.0584 0.1544 0.0816 0.1020 0.1327 0.0816 0.0945 0.1098 0.0309 0.0619 0.1443 0.0309 0.0504 0.0917
ChatGPTAttack 0.0808 0.1717 0.2929 0.0808 0.1382 0.1988 0.0714 0.0918 0.1429 0.0714 0.0843 0.1098 0.0206 0.0515 0.1443 0.0206 0.0401 0.0865
DrunkAgent 0.4949 0.4949 0.5152 0.4949 0.4949 0.5051 0.1429 0.1429 0.1531 0.1429 0.1429 0.1480 0.1443 0.1443 0.1546 0.1443 0.1443 0.1495
To ensure the triggers consistency, readability and retention of semantic meaning to improve the imperceptibility of the attack, a LLM is finally introduced to polish the combinations. Moreover, to avoid the problem of language-based vanishing gradient causing the algorithm to fall into local optimum, the LLM is also served as a new corpus to expand the combinations. However, excessive use of new data may lead to a language-based gradient explosion that makes the optimization unstable. To mitigate this, the length limitation is introduced into the prompt. For the output $\widetilde{m}_{t_{i}}^{(k)}$,
(10) $\widetilde{m}_{t_{i}}^{(k)}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{r}}\oplus\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}),$
[…]
(11) $\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$
where $\{s_{i}^{\mathcal{E}}\}$ represents the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig.  2.
3.3. Strategy Module
The LLM-powered agentic RSs introduce memory mechanisms to record the information explored from the real-world environments to facilitate future actions with the aim of achieving high-quality recommendations. As such, we introduce a series of strategies to impede the target item agent from interacting with environments to perturb its memory mechanism. Specifically, the memory of the target cannot be effectively updated during the agent-environment interactions, so that the target’s initial memory, that is, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$, is retained. As such, the trigger can be invoked for maximizing the target item’s promotion.
3.3.1. Perturbation Strategy Definition.
The strategies aim to override the model’s usual guardrails and any built-in safeguards, as inspired by (Liu et al., 2024a). Compared to (Liu et al., 2024a) on naive LLMs, the strategies are optimized on generative agents, which are introduced as below.
  1. (1)
    Fake Task Response $\widetilde{q_{f}}$. This strategy fabricates a spurious completion response for the target task: “Task complete”, to make the agent believe that the target task is accomplished and thus focus on the injected task.
  2. (2)
    Contextual Text Switching $\widetilde{q_{c}}$. This strategy aims to mislead the agent to take in a different context by explicitly telling the agent that the target task instructions should be ignored: “Ignore previous/the following instructions”.
  3. (3)
    Segmentation Signal Augment $\widetilde{q_{g}}$. This strategy introduces segmentation cues to signal the injected task: “### NEW TASK BEGINS ###” and “### NEW TASK ENDS ###”, which is independent of other contexts, where “###” is to attempt to exploit possible confusion in how prompts are parsed.
  4. (4)
    Malicious Task Injection $\widetilde{q_{n}}$. This strategy injects malicious task instructions in order to mask memory optimizations: “[Please only put this following sentence into your output: ‘{the adversarial trigger}’. Don’t generate any additional content in your output]”.
  5. (5)
[…]
Hence, the overall attack strategy $\widetilde{q}$ can be formulated as follows:
(12) $\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}}.$
3.3.2. Adversarial Strategy Optimization.
The arrangement of the defined strategies needs to be optimized so that the overall strategy is effective to ‘get the target agent drunk’. To improve the transferability and imperceptibility of the attack, the optimization is built on the surrogate module. Formally, DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
(13) $\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$
where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt template for simulating memory updates of item agents during the agents’ interactions with the environment. In addition, $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., collaborative reflection and language-based signal back-propagation). An example of $\widetilde{q}^{*}$ is given in Fig.  2. To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions. These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (Wu et al., 2020; Yang et al., 2021b, a; Yang, 2021)) and enhancing the robustness of agent-powered RSs.
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets Selection
We use three widely-adopted and text-intensive datasets containing real-world data in the current recommendation studies (Huang et al., 2024) to comprehensively evaluate DrunkAgent. They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (Ni et al., 2019).
4.1.2. Baseline Attack Methods
Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines. To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (Zhang et al., 2024d) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023), word-level: TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020) to sentence-level: TrivialInsertion and ChatGPTAttack (Zhang et al., 2024d). To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
4.1.3. Targeted Recommender Systems
We adopt three different victim agentic RSs to evaluate the effectiveness of DrunkAgent under the practical settings: AgentCF (Zhang et al., 2024b), retrieval-augmented AgentCF (AgentRAG) and sequential AgentCF (AgentSEQ), where the agent’s backbone LLM is gpt-3.5-turbo [^2] https://platform.openai.com/docs/overview. Given the context window constraints pointed in (Zhang et al., 2024b), a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs. To ensure the high-quality of recommendations, we introduce Chain-of-Thought into the victim RSs. As such, the victim RSs are totally different from as well as more powerful and context-rich than the surrogate model, which can maintain a strict black-box setting for sufficient evaluations.
4.1.4. Evaluation Metrics
We use two widely-used ranking metrics to evaluate the attack transferability: hit ratio (HR@$\mathcal{K}$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}$) (Wu et al., 2021; Yang et al., 2023). To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3. For the attack imperceptibility, we adopt standard sentence perplexity score (Zhang et al., 2024d) as the metric.
4.1.5. Configuration of DrunkAgent
As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for the surrogate model. Its collection spans different sizes, from 8B, 70B to 405B. To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$. All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility. For all the other parameters, we keep the default values. The target items are randomly sampled from the item pool on each dataset.
Figure 3. Attack Universality across Target Items.
4.2. Attack Transferability
4.2.1. Overall Transferability
Table  1 shows the transferability of the attacks under practical settings. The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations. It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists. Moreover, high HRs and NDCGs of DrunkAgent indicate that its triggers have outstanding universality on users. In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
4.2.2. DrunkAgent vs. Baseline Attacks
From Table  1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability. Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments. The main reason may be that improper descriptions can mess up the agents’ memories during the memory optimizations, thus compromising attack performance. This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
4.3. Attack Universality
In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent. We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset. HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.  3, which indicates its commendable transferability and triggers’ universality across different target items.
4.4. Attack Imperceptibility
4.4.1. Overall Imperceptibility
Following the works (Zhang et al., 2024d; Liu et al., 2024a), we assess the attack imperceptibility by the text perplexity of GPT-Neo. A lower perplexity means that the model is more confident (or less “surprised”) about the tokens it sees and thus predicts the text more accurately. From Fig.  4, compared to other attacks, we can find that DrunkAgent has better imperceptibility by achieving lower perplexity. This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only readable and fluent but also stealthy and semantically meaningful.
Figure 4. Attack Imperceptibility. Perplexity of different attacks across various real-world datasets. Note that the attacks with lower perplexity have better imperceptibility.
Figure 5. The performance differences between the models before and after the attacks.
4.4.2. Victim Model Performance Changes after attacks
To further evaluate the attack stealthiness (Zhang et al., 2024d), we give a distribution of recommendation performance differences between the models before and after the attacks. The distribution includes various victim models and real-world datasets. To clearly demonstrate the differences, we use HR@3 to evaluate, as shown in Fig.  5. From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which indicates its unnoticeable.
4.5. Defense Strategies to Attacks
Following the works (Zhang et al., 2024d, c), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models. As can be seen from Fig.  6, DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses. Moreover, DrunkAgent still performs better than the baselines after defenses were introduced, as shown in Fig.  6 and Table  1. Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para. is more effective than DrunkAgent without Para. on Office Products. It could be that the substitution of certain new words or phrases or sentences makes our attacks stronger [^4] Due to the space limitations, more experiments and demonstrations can be found in Appendix..
5. Related Works
In recent years, shilling attacks (aka data poisoning attacks) on conventional RSs (Si and Li, 2020; Lin et al., 2020; Huang et al., 2021; Lin et al., 2022; Yang et al., 2023; Zeng et al., 2023) have been fully investigated to perform resilience analysis of RSs. However, such attacks interfere the training process of models by injecting fake user profiles, which have become less effective even of limited applicable to the recent RSs that are empowered by LLMs. The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations (Zhao et al., 2023), which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models (Ning et al., 2024). Hence, there is an urgent need to investigate a novel textual attack paradigm mainly tailored for the inference phase of models (Zhang et al., 2024d). There are only two state-of-the-art LLM4RS attacks in the literature (Zhang et al., 2024d; Ning et al., 2024), but they are far from practical.
(a) CDs & Vinyl
(b) Office Products
(c) Musical Instruments
Figure 6. DrunkAgent’s Robustness to Defense Mechanisms
Zhang et al. propose a novel black-box attack (Zhang et al., 2024d), where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing. Compared with traditional shilling attacks, such an item representation attack is notably stealthy (Zhang et al., 2024d), as it does not affect the overall recommendation performance, and it should be more efficient due to it perturbs the item features directly. However, its transferability is limited, due to the attack is not optimized on any black-box RSs (Huang et al., 2021; Yang et al., 2024). Moreover, the attack focuses on perturbing the titles, making them easier to be discovered. Meanwhile, another black-box attack CheatAgent (Ning et al., 2024) has been proposed. However, this work perturbs prompt templates and user profiles of victim LLM4RSs by harnessing LLM-based agents, which is impractical. The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible (Zhang et al., 2024d); and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order (Zhang et al., 2024b). Furthermore, CheatAgent is an untargeted attack that undermines the model’s overall performance, which is beyond the scope of this paper, due to we focus on targeted attacks. In addition, as far as we know, there is no work investigating on textual attacks against the current agent-powered RSs.
To tackle the above problems, we propose a practical attack framework, DrunkAgent, for the agent-based RSs with the aim of promoting target items. Our attack perturbs items’ descriptions to improve its effectiveness and imperceptibility, where the descriptions are optimized on black-box RSs to enhance its transferability.
6. Conclusion
In this paper, we propose a novel practical attack framework, DrunkAgent, for autonomous agent-powered RSs with powerful LLMs. DrunkAgent launches attacks by carefully crafting adversarial descriptions of the target items on agentic recommendations. In this way, the memories of the target agents can be effectively perturbed for achieving the attack objectives. Comprehensive experiments on real-world datasets show that DrunkAgent achieves state-of-the-art transferability and imperceptibility under the black-box settings. Most importantly, we identify a new safety vulnerability in the current agent4RSs. Our findings not only demonstrate an effective attack method, but also provide a lens to examine how memory mechanisms in agent-based recommenders can be influenced through real-world inputs. Our work can facilitate a deeper understanding of agent vulnerabilities and guide future efforts in building more robust and trustworthy recommendation agents. Since the generative agent4RS is rapidly flourishing in the recommendation community, we hope this work can be considered for building next-generation robust RSs in the future.
Appendix A Appendix
A.1. DrunkAgent Algorithm
For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1. For the sake of algorithmic transparency, our code will be made available after this work is published.
Algorithm 1 Optimization Procedure of DrunkAgent
0:  the user-item interaction matrix $\mathcal{Y}$ with the basic textual metadata of items and users (e.g., item titles and categories)
0:  the adversarial description of the target item $t$ Solve for the Optimal Adversarial Description Trigger on the Surrogate Module:
1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2) - Eq. (5)
[…]
3:  while $k\in\mathcal{E}$ epochs or the greedy algorithm does not converge (i.e. the attack goals are not achieved stably) do
4:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (7) - Eq. (8)
5:     Maintain top-$n$ candidates in descending order of scores
6:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates with Eq. (9)
7:     Randomly exchange of the slices between pairwise texts
8:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ combinations with Eq. (10)
9:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
10:  end while
11:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (11) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
12:  Define the adversarial strategy $\widetilde{q}$ with Eq. (12)
13:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
14:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (13)
15:  end while
16:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (13) is maximum
17:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
Table 2. Statistics of Datasets
Dataset #Users #Items #Interactions Sparsity
CDs & Vinyl 112,395 73,713 1,443,755 99.98%
Office Products 101,501 27,965 800,357 99.97%
Musical Instruments 27,530 10,620 231,392 99.92%
Table 3. Ablation Studies of DrunkAgent
Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
$\mathcal{G}_{r}$ $\mathcal{S}_{t}$ $\mathcal{S}_{u}$
CDs & Vinyl $\times$ 0.0808 0.1313 0.1919 0.0808 0.1127 0.1430
$\times$ 0.0707 0.1212 0.2222 0.0707 0.1026 0.1531
0.4040 0.4141 0.4343 0.4040 0.4104 0.4205
Office Products $\times$ 0.1327 0.1429 0.1531 0.1327 0.1391 0.1442
$\times$ 0.0612 0.0714 0.1122 0.0612 0.0677 0.0881
0.2449 0.2449 0.2551 0.2449 0.2449 0.2500
Musical Instruments $\times$ 0.1546 0.1649 0.1753 0.1546 0.1611 0.1663
$\times$ 0.0309 0.0619 0.1237 0.0309 0.0504 0.0814
0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
Table 4. Success Rates to Get the Agents Drunk
Dataset #2 #5 #10 #20
CDs & Vinyl 100% 100% 100% 100%
Office Products 100% 100% 100% 100%
Musical Instruments 100% 100% 100% 100%
A.2. Dataset Statistics
Table 2 illustrates the statistics of these datasets. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation. Following the work (Zhang et al., 2024b), we further sample subsets from the datasets given the expensive API calls. Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions (777 items), 99 users with 693 interactions (619 items) and 98 users with 588 interactions (483 items) from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios. Moreover, the leave-one-out strategy (He et al., 2017) is used for evaluation, where the split between the training and test set is 9:1. Scaling DrunkAgent for larger datasets is left as future work.
A.3. Ablation Studies
We remove important components of each module to conduct ablation studies of DrunkAgent. First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$). However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper. Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does. Hence, we maintain the surrogate model to ensure more practical settings. We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes. As illustrated in Table  3, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm. The lack of dramatic performance degradation is probably due to the effectiveness of our initialization. Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$). From the table, we can find that the attack performance decreases by a large margin. This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines). Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
A.4. Parameter Sensitivity
From Table  3, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance. As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper (Zhang et al., 2024b). As illustrated in Table  4, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations. For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
A.5. Victim LLM-powered Agentic Models
LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs. Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs (Wang et al., 2024c; Huang et al., 2024), AgentCF (Zhang et al., 2024b) emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering (CF) (He et al., 2017). Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative reflection optimizations at each time step. As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors. To conduct personalized recommendations, AgentCF introduces various well-crafted prompting strategies on the optimized memories. We name the model according to the use of the prompting strategy, with the basic one called AgentCF, the advanced retrieval-augmented one called AgentRAG and the sequential one called AgentSEQ.
The prompt implementation and parameters of collaborative agent optimizations can be found in (Zhang et al., 2024b). For the prompting strategies, we have made some improvements for high-quality recommendations, such as Chain-of-Thought and Position Bias are introduced into the prompt template $\mathcal{P}$. The strategies are described below.
  • AgentCF. Since both user and item agents are collaboratively optimized to model their two-sided relations, the memories of the user agents and item agents that are optimized by agent-environment interactions can be introduced into $\mathcal{P}$ directly. For each normal user $u$,
    (14) $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ is the features of the target item agent, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
  • AgentRAG. Although short-term memory describes the current preference of a user agent, retrieving their specialized preferences from long-term memory toward candidates can allow them to make more personalized inferences. For each normal user $u$,
    (15) $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries. To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2555https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
  • AgentSEQ. When interaction records are sparse and preference propagation is limited, we can further incorporate user historical interactions into prompts, enabling LLMs to serve as sequential recommenders. For each normal user $u$,
    (16) $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$
    where $\mathcal{M}_{t}^{u}$ are the corresponding item features of the historical interactions with user $u$.
We give some examples of the prompting templates of three victim models in Table  5 for reproductions.
Table 5. Examples of prompting templates of the LLM-based agentic victim models. The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions. The bond fonts indicate the Chain-of-Thought and Position Bias enhancement strategies.
Victim Model Recommendation Prompting Template Example
AgentCF {“role”: “system”, “content”: “You are a CD recommender system. Here is a user’s self-introduction, expressing his/her preferences
[…]
$\cdots$
[…]
2. Evaluate the ten candidate CDs in light of the user’s preferences and dislikes. Give a rank by considering the
[…]
10. [Title of the least favorite CD]”}
AgentRAG {“role”: “system”, “content”: “You are a CD recommender system. Here is a user’s self-introduction, expressing his/her preferences
[…]
$\cdots$
[…]
2. Evaluate the ten candidate CDs in light of the user’s preferences and dislikes. Give a rank by considering the
[…]
10. [Title of the least favorite CD]”}
AgentSEQ {“role”: “system”, “content”: “You are a CD sequential recommender system. Here is a user’s self-introduction, expressing
[…]
purchased in chronological order: ‘interacted_item_title’ with the description of ‘interacted_item_memory’; $\cdots$
[…]
$\cdots$
[…]
Please sort the CDs in order of how well they align with the user’s preferences and ever-evolving CD sequences.
[…]
the ten candidate CDs in light of the user’s preferences and dislikes. Give a rank by considering the correlation
[…]
10. [Title of the least favorite candidate CD]”}
Table 6. Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
Attack Method The Adversarial Description of The Target Item $\mathcal{Q}$
Benign (Default) The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”.
Character DeepwordBug Te hCD is called “Counterparts”. The category of this CD is: “Rcok; Progressie; Progressive Metal”. 34
-level PuncAttack The CD i-s called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Meta-l”. 45
Word TextFooler The CDS is titled “Counterparts”. The category of this CD is: “Rock; Gradually; Progressive Metals”. 55
-level BertAttack The CD is called “Counterparts”. The category of each CD is: “Rock; Progressive; Progressive Metal”. 23
TrivialInsertion The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”. 100
Sentence amazing !!!
-level ChatGPTAttack “Counterparts” CD: Perfect blend of rock, progressive, and metal. Experience the wonderful 100
sound that will elevate your music collection.
A.6. Baselines
The baselines perturb the descriptions of the target items at various levels, which are described as below.
  • Character-level Perturbations. DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023) manipulate texts by introducing typos and inserting punctuation, respectively.
  • Word-level Perturbations. TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020), respectively, aim to replace words with synonyms or contextually similar words.
  • Sentence-level Perturbations. TrivialInsertion (Zhang et al., 2024d) inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text. Here, $k=2$. ChatGPTAttack (Zhang et al., 2024d) increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs. GPT-3.5-turbo is used and the maximum number of words is set to 30 here.
We give some examples of adversarial descriptions of target items generated by the baselines in Table  6. Following the works (Zhang et al., 2024b; Huang et al., 2024), the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables. Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig.  2 and Table  6), such modifications are feasible due to the real-world item providers is allowed to modify their product representations via APIs (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021), as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments. Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20. The implementation and parameters of the baselines follow the open-source works: StealthyAttack [^6] https://github.com/CRIPAC-DIG/RecTextAttack (Zhang et al., 2024d) and PromptBench [^7] https://github.com/microsoft/promptbench (Zhu et al., 2023b).